OpenStudy (anonymous):

Ben joined a book club. The company sent him 12 books of his choice to welcome him to the club. Then he received a regular number of books each month. After 18 months, he had a collection of 54 books. Which equation can be used to calculate the number of books Ben will have at any number of months after joining the club?

OpenStudy (rulnick):

Well, he starts with 12 and then gets n per month. So if the number of months so far is m, the total number of books received would be 12 + mn right?

OpenStudy (rulnick):

For example, after 18 months: 12 + 18n where, again, n is the number of books each month.

OpenStudy (anonymous):

so it would be b = 54 - 12m

OpenStudy (rulnick):

No, I think it would be 54 = 12 + 18n, and then solve for n (or b, if you prefer b instead of n: 54 = 12 + 18b). Are you sure the numbers are right? It wouldn't work out to a whole number for n (or b).